OpenStudy (anonymous):

how do I solve for x? 2^(2x+3)= 32

OpenStudy (anonymous):

32 = 2^5 Therefor: 2^(2x+3)=2^5 ==> 2x+3 = 5 ==> 2x = 2 ==> x = 1 To check your answer: 2^(2*1+3) = 2^5 = 32 Done

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

You're welcome